OpenStudy (nuccioreggie):

A scale drawing of Jerome's living room is shown below: A rectangle is shown. The length of the rectangle is labeled as length equal to 8 cm, and the width is labeled as width equal to 6 cm. If each 2 cm on the scale drawing equals 6 feet, what are the actual dimensions of the room? Length = 16 feet, width = 12 feet Length = 11 feet, width = 9 feet Length = 24 feet, width = 18 feet Length = 20 feet, width = 18 feet

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

we have to compute this: length = (8/2)*6=...feet width=(6/2)*6=...feet

OpenStudy (nuccioreggie):

Its D right

OpenStudy (michele_laino):

are you sure?

OpenStudy (nuccioreggie):

Am i right

OpenStudy (michele_laino):

no, I'm sorry

OpenStudy (nuccioreggie):

What is the correct choice

OpenStudy (michele_laino):

what is (8/2) *6=...?

OpenStudy (nuccioreggie):

24 so its C

OpenStudy (michele_laino):

that's right!